OpenStudy (anonymous):

Please! help me check my answe,do even only, 2) Degree 4) GCF 6) Polynomial 8) Simplify 10) Synthetisc dubtitution

OpenStudy (anonymous):

sorry subtitution not dubtitution

OpenStudy (radar):

8 = 8 10=4

OpenStudy (anonymous):

not solve, just fill the word in the blank

OpenStudy (anonymous):

I need check my answer if I correct

OpenStudy (radar):

8, Simplify 4, synthetic division

OpenStudy (radar):

seems the choices are too restricted!!

OpenStudy (anonymous):

number 10 Synthetisc dubtitution ?

OpenStudy (radar):

Not for sure, have you googled those words?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have time look another problem?

OpenStudy (radar):

I don't see any on the list that would match "degree"

OpenStudy (radar):

I may be of help, what is the probllem?

OpenStudy (anonymous):

you want me sent you a lesion degree?

OpenStudy (radar):

the degree of an equation depends on the highest exponent of the variable, and there was nothing like that on the list. What is the next problem?

OpenStudy (radar):

I have looked it over. Here is what I have come up with: Let x = longer leg Let y= shorter leg then 24-(x+y)= hypotenuse You with me so far. Ask question now about what I have set up for solving this problem.

OpenStudy (radar):

We are told the following: 3x-2y=24-(x+y)+3 Do you see where that come from?

OpenStudy (anonymous):

P: x+y+z=24 The given: (3*x)-(2*y)=h+2 I know from here ,But I don't know next

OpenStudy (anonymous):

sorry z for hypotenuse The given: (3*x)-(2*y)=z+2

OpenStudy (radar):

O.k I put 3 where it should be 2, my equation should be 3x-2y = 24-x-y +2

OpenStudy (radar):

then we would have 4x+3y = 26 That is an equation with two unknowns we need to make another equartion. But first need to fix another mistake. 4x-y=26 Do you see why 4x-y=26 cm????

OpenStudy (anonymous):

yes I see

OpenStudy (radar):

This means y=4x-26. Problem says it is a right triangle we know that by pythagorem theorem hypotenuse is equal to the square root of the sum of the legs squared or\[\sqrt{x ^{2}+y ^{2}}\]

OpenStudy (radar):

we substitute for y and get this: \[hyp=\sqrt{x ^{2}+(4x-26)^{2}}\] We also know that the hypotenuse is equal to 24-x-y substituting for y we now have hypotnuse=24-x-(4x-26) hyp=-5x+50 \[-5x+50=\sqrt{x ^{2}+(4x-26)^{2}}\]Now we have an equality with one unknown

OpenStudy (radar):

The radical should extend completely across\[-5x+50=\sqrt{x ^{2}-(16x ^{2}-208x+676}\]

OpenStudy (anonymous):

Arte you still there?

OpenStudy (anonymous):

From here I just solve for x?

OpenStudy (radar):

square both sides of the equation \[25x ^{2}-500x+2500=x ^{2}-16x ^{2}-208x+676\] This thing keeps getting bigger and bigger!!!\[40x ^{2}-292x+1824=0\] A quadratic that should be solvable lol

OpenStudy (anonymous):

Ok. thank I can do from now

OpenStudy (radar):

You can divide by 4 and get \[10x ^{2}-73x+456=0\]

OpenStudy (anonymous):

I have one more very quick, can you show me how to write faction on the bottom of the chat, I keep play but I can do it.

OpenStudy (radar):

I don't know how to write frac on the equation line, if you google latex you will find the math languate they use here.

OpenStudy (anonymous):

on the bottom of the chat window

OpenStudy (anonymous):

it say Equation

OpenStudy (anonymous):

I see every people write fraction

OpenStudy (radar):

loki showed me how, but I forget, if you look at "latex" it will explain it.

OpenStudy (anonymous):

it Ok I will ask pepole around, thank you for spend time solve my problem

OpenStudy (anonymous):

I have to go now bye

OpenStudy (radar):

Ok, somewhere there is an error in the solution as the root is a complex value. I will loolk it over and see where it went wrong.

OpenStudy (radar):

There were errors, the final quadratic should appear as:\[8x ^{2}-292x+1824=0\] When solved, you will find x=8 (longer leg) substitute x in the y=4x-26=32-26=y y=6 cm (shorter leg) 24-6-8=10cm the hypotenuse It checks out.